OpenStudy (anonymous):

Which equation is equivalent to 4(2−5x)=6−3(1−3x) ? A 8x = 5 B 8x =17 C 29x = 5 D 29x =17

OpenStudy (radar):

8-20x=6-3+9x -29x=-5 29x=5 I would go with C

OpenStudy (anonymous):

How does 6-3+9x = -5?

OpenStudy (anonymous):

I also don't understand how 8-20x = -29x. If you could draw out how you solved the equation or expanded a little on the steps part I'd really appreciate it.

OpenStudy (radar):

A better question is how does 6-3-8 = -5 You are subtracting 8 from both sides, noticed it is no longer on the left.

OpenStudy (anonymous):

Ah, I see..

OpenStudy (radar):

Using algebra you may add or subtract from one side as long as you treat both sides the same way, I also subtracted 9x from both sides.

OpenStudy (radar):

The next time I will show it more clearly. Do you follow with understanding?

OpenStudy (anonymous):

Yes, I know how you got the answer now. Thank you very much :)

OpenStudy (radar):

You're welcome, and good luck with those problems.